OpenStudy (anonymous):

what is the smallest number of portriats that can be distributed equally among either four grandparents and six aunts and uncles with no portraits left over?

OpenStudy (anonymous):

you need to find the common denominator of 6 and 4

OpenStudy (anonymous):

if you had this problem \[\frac{1}{4}+\frac{1}{6}\] what would you do

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i guess add it 2 over ten .or 2over five not sure dont understand

OpenStudy (anonymous):

i guess its 2

OpenStudy (anonymous):

the common denominator

OpenStudy (anonymous):

not quite the common denominator is going to be a number of a factor or 4 or 6 so go through each factor 1*4=4 2*4=8 3*4=12 4*4=16 now do 6 1*6=6 2*6=12 3*6=18 you can see that both have 12, therefore 12 is your common denominator in which 4 and 6 go evenly into

OpenStudy (anonymous):

so your answer is 12 if you do this the long way you can see.. first you need 6 pictures because the second part of the problem distributes them to 6 people for 6, it is evenly distributed for 6 but you're left over with 2 for 4 people for 7, you have 3 leftover for 4 people and 1 leftover for 6 people for 8, you have none leftover for 4 people but 2 leftover for 6 people for 9, you have 1 leftover for 4 people, and 3 for 6 people for 10, you have 2 leftover for 4 people and 4 for 6 people for 11, you have 3 leftover for 4 people and 5 leftover for 6 people 12 you give 3 pictures to the each person for the 4 people and 2 to each for the 6 peopl

OpenStudy (anonymous):

thanks alot ive never been good at math at all thanks bunches

OpenStudy (anonymous):

Look at the pattern shown below. What number is missing from the pattern? 20, 22, 26, 34, __,8